OpenStudy (kitkat1):

Celine has a bottle that contains 20% of milk and the rest water. The bottle has 1 liter of water. Part A: Write an equation in one variable that can be used to find the total number of liters of milk and water in the bottle. Define the variable used in the equation. Part B: How many liters of milk are present in the bottle? Show your work.

OpenStudy (kitkat1):

I will give medals please help me

OpenStudy (anonymous):

for B it 1/4 liter

OpenStudy (kitkat1):

How did u get that?

OpenStudy (anonymous):

it state that 20% of it was milk and the rest is water 100-20=80

OpenStudy (anonymous):

so 80% of the bottle is 1 liter

OpenStudy (anonymous):

divide it by 4 it become 20%=1/4 liter

OpenStudy (kitkat1):

Okay but do u know the first part?

OpenStudy (anonymous):

just maybe

OpenStudy (anonymous):

\[\frac{ 1 }{ 5 }x +1=x\]

OpenStudy (kitkat1):

Okay thanks